# Carrier Aggregation Enabled MIMO-OFDM Integrated Sensing and Communication

### v1

@@ VI Conclusion
In this paper, CA-enabled MIMO-OFDM ISAC system is investigated, where CA aggregates fragmented high and low-frequency bands, achieving the high-speed communication and high-accuracy sensing. Firstly, we introduce the signal model of the proposed CA-enabled MIMO-OFDM ISAC system and the ISAC signal processing framework. To address the challenges in the sensing signal processing introduced by CA, symbol-level fusion method of the sensing data on high and low-frequency bands is proposed. In the proposed symbol-level fusion method, the phase alignment of the echo signals on high and low-frequency bands is realized by the angle compensation, space-domain diversity and vector cross-correlation operations. Then, the fusion algorithms of feature vectors are proposed to realize high-accuracy sensing. The simulation results reveal that the performance of symbol-level fusion method is better than that of data-level fusion method. Furthermore, the closed-form communication MI and sensing CRLB of the proposed ISAC signal are derived. Numerical results reveal that the proposed CA-enabled MIMO-OFDM ISAC signal exhibits superior communication rate and high-accuracy sensing performance over the conventional MIMO-OFDM ISAC signal. This work reveals the potential of CA-enabled MIMO-OFDM ISAC signal in enhancing communication and sensing capabilities.
@@
According to (12), the received echo signal by the $p$-th received antenna on the $n$-th subcarrier during the $m$-th OFDM symbol time is expressed as
$y_{p,m,n}=$ $Ae^{j2\pi mf_{\mathrm{C}}^{b}T^{b}\gamma}e^{-j2\pi n\Delta f^{b}\tau_{0}}e^{j2\pi p\left(\frac{d_{\text{r}}}{\lambda^{b}}\right){\sin\theta_{\text{Rx}}}}$ (66)
$+w_{p,m,n},$
where $A\in\mathbb{C}$ and $w_{p,m,n}\sim\mathcal{CN}(0,\sigma^{2})$ is the AWGN.
The likelihood function for the joint estimation of $\sin\theta_{\text{Rx}}$, $\tau_{0}$, and $\gamma$ is
$\ln f\left(y;\sin\theta_{\text{Rx}},\tau_{0},\gamma\right)$ (67)
$=-\frac{M_{b}N_{b}N_{\text{U}}}{2}\ln\left(2\pi\sigma^{2}\right)$
$\quad-\frac{1}{2\sigma^{2}}\sum_{p}\sum_{m}\sum_{n}\left(y_{p,m,n}-s_{p,m,n}\right)^{*}\left(y_{p,m,n}-s_{p,m,n}\right),$
where the PDF is
$f\left(y;\sin\theta_{\text{Rx}},\tau_{0},\gamma\right)$ (68)
$=\frac{1}{\left(2\pi\sigma^{2}\right)^{M_{b}N_{b}N_{\text{U}}/2}}e^{-\frac{1}{2\sigma^{2}}\sum\limits_{p}\sum\limits_{m}\sum\limits_{n}\left|y_{p,m,n}-s_{p,m,n}\right|^{2}},$
@@
$s_{p,m,n}=Ae^{j2\pi mf_{\mathrm{C}}^{b}T^{b}\gamma}e^{-j2\pi n\Delta f^{b}\tau_{0}}e^{j2\pi p\left(\frac{d_{\text{r}}}{\lambda^{b}}\right){\sin\theta_{\text{Rx}}}}.$ (69)
@@
$\mathbf{F}^{-1}$ $=\left[\begin{array}[]{ccc}F_{\sin{\theta_{\text{Rx}}},\sin{\theta_{\text{Rx}}}}&F_{\sin{\theta_{\text{Rx}}},\tau_{0}}&F_{\sin{\theta_{\text{Rx}}},\gamma}\\ F_{\tau_{0},\sin{\theta_{\text{Rx}}}}&F_{\tau_{0},\tau_{0}}&F_{\tau_{0},\gamma}\\ F_{\gamma,\sin{\theta_{\text{Rx}}}}&F_{\gamma,\tau_{0}}&F_{\gamma,\gamma}\end{array}\right]^{-1}$ (70)
$=\left[\begin{array}[]{ccc}CRLB(\sin{\theta_{\text{Rx}}})&Cov(\sin{\theta_{\text{Rx}}},\tau_{0})&Cov(\sin{\theta_{\text{Rx}}},\gamma)\\ Cov(\tau_{0},\sin{\theta_{\text{Rx}}})&CRLB(\tau_{0})&Cov(\tau_{0},\gamma)\\ Cov(\gamma,\sin{\theta_{\text{Rx}}})&Cov(\gamma,\tau_{0})&CRLB(\gamma)\end{array}\right],$

### v2

@@ VI Conclusion
In this paper, CA-enabled MIMO-OFDM ISAC system is investigated, where CA aggregates fragmented high and low-frequency bands, achieving the high-speed communication and high-accuracy sensing. Firstly, we introduce the signal model of the proposed CA-enabled MIMO-OFDM ISAC system and the ISAC signal processing framework. To address the challenges in the sensing signal processing introduced by CA, symbol-level fusion method of the sensing data on high and low-frequency bands is proposed. In the proposed symbol-level fusion method, the phase alignment of the echo signals on high and low-frequency bands is realized by the angle compensation, spatial filtering and cyclic cross-correlation operations. Then, the fusion algorithms of feature vectors are proposed to realize high-accuracy sensing. The simulation results reveal that the performance of symbol-level fusion method is better than that of data-level fusion method. Furthermore, the closed-form communication MI and sensing CRLB of the proposed ISAC signal are derived. Numerical results reveal that the proposed CA-enabled MIMO-OFDM ISAC signal exhibits superior communication rate and high-accuracy sensing performance over the conventional MIMO-OFDM ISAC signal. This work reveals the potential of CA-enabled MIMO-OFDM ISAC signal in enhancing communication and sensing capabilities. However, the proposed fusion method does not reach the theoretical lower bound of the proposed CA-enabled MIMO-OFDM ISAC system. Future trends include designing fusion methods and resource allocation methods to approximate the theoretical lower bound.
@@
According to (12), in the $b$-th CC, the received echo signal of the $i$-th target by the $p$-th receive antenna on the $n$-th subcarrier during the $m$-th OFDM symbol time is expressed as
$y_{p,m,n}^{i}=$ $Ae^{j2\pi mf_{\mathrm{C}}^{b}T^{b}\gamma_{i}}e^{-j2\pi n\Delta f^{b}\tau_{i,0}}e^{j2\pi p\left(\frac{d_{\text{r}}}{\lambda^{b}}\right){\sin\theta_{i,\text{Rx}}}}$ (64)
$+w_{p,m,n}^{i},$
where $A\in\mathbb{C}$ and $w_{p,m,n}^{i}\sim\mathcal{CN}(0,\sigma^{2})$ is the AWGN.
The likelihood function for the joint estimation of $\sin\theta_{i,\text{Rx}}$, $\tau_{i,0}$, and $\gamma_{i}$ is
$\ln f\left(y;\sin\theta_{i,\text{Rx}},\tau_{i,0},\gamma_{i}\right)$ (65)
$=-\frac{M_{b}N_{b}N_{\text{R}}}{2}\ln\left(2\pi\sigma^{2}\right)$
$\quad-\frac{1}{2\sigma^{2}}\sum_{p}\sum_{m}\sum_{n}\left(y_{p,m,n}^{i}-s_{p,m,n}^{i}\right)^{*}\left(y_{p,m,n}^{i}-s_{p,m,n}^{i}\right),$
where $y$ is the received echo signal of the $i$-th target, and the PDF is
$f\left(y;\sin\theta_{i,\text{Rx}},\tau_{i,0},\gamma_{i}\right)$ (66)
$=\frac{1}{\left(2\pi\sigma^{2}\right)^{M_{b}N_{b}N_{\text{R}}/2}}e^{-\frac{1}{2\sigma^{2}}\sum\limits_{p}\sum\limits_{m}\sum\limits_{n}\left|y_{p,m,n}^{i}-s_{p,m,n}^{i}\right|^{2}},$
@@
$s_{p,m,n}^{i}=Ae^{j2\pi mf_{\mathrm{C}}^{b}T^{b}\gamma_{i}}e^{-j2\pi n\Delta f^{b}\tau_{i,0}}e^{j2\pi p\left(\frac{d_{\text{r}}}{\lambda^{b}}\right){\sin\theta_{i,\text{Rx}}}}.$ (67)
@@
$\mathbf{F}^{-1}=\left[\begin{array}[]{ccc}F_{\sin{\theta_{i,\text{Rx}}},\sin{\theta_{i,\text{Rx}}}}&F_{\sin{\theta_{i,\text{Rx}}},\tau_{i,0}}&F_{\sin{\theta_{i,\text{Rx}}},\gamma_{i}}\\ F_{\tau_{i,0},\sin{\theta_{i,\text{Rx}}}}&F_{\tau_{i,0},\tau_{i,0}}&F_{\tau_{i,0},\gamma_{i}}\\ F_{\gamma_{i},\sin{\theta_{\text{Rx}}}}&F_{\gamma_{i},\tau_{i,0}}&F_{\gamma_{i},\gamma_{i}}\end{array}\right]^{-1}$ (68)
$=\left[\begin{array}[]{ccc}CRLB(\sin{\theta_{i,\text{Rx}}})&Cov(\sin{\theta_{i,\text{Rx}}},\tau_{i,0})&Cov(\sin{\theta_{i,\text{Rx}}},\gamma_{i})\\ Cov(\tau_{i,0},\sin{\theta_{i,\text{Rx}}})&CRLB(\tau_{i,0})&Cov(\tau_{i,0},\gamma_{i})\\ Cov(\gamma_{i},\sin{\theta_{i,\text{Rx}}})&Cov(\gamma_{i},\tau_{i,0})&CRLB(\gamma_{i})\end{array}\right],$